### v1

# AoA Services in 5G Networks: A Framework for Real-World Implementation and Systematic Testing

Abstract
Accurate positioning is a key enabler for emerging 5G applications. While the standardized Location Management Function (LMF) operates centrally within the core network, its scalability and latency limitations hinder low-latency and fine-grained localization. A practical alternative is to shift positioning intelligence toward the radio access network (RAN), where uplink sounding reference signal (SRS)-based angle-of-arrival (AoA) estimation offers a lightweight, network-native solution. In this work, we present the first fully open-source 5G testbed for AoA estimation, enabling systematic and repeatable experimentation under realistic yet controllable channel conditions. The framework integrates the NVIDIA Sionna RT with a Keysight PROPSIM channel emulator and includes a novel phase calibration procedure for USRP N310 devices. Experimental results show sub-degree to few-degree accuracy, validating the feasibility of lightweight, single-anchor, network-native localization within next-generation 5G systems.
I Introduction
The rapid evolution of wireless communication systems has brought about new demands not only in terms of data throughput and latency but also in high-precision positioning, especially for use cases such as autonomous vehicles, drone navigation, and Internet of Things (IoT) networks [1]. Traditional positioning techniques, such as Global Navigation Satellite System (GNSS), often suffer from limited coverage or poor accuracy in indoor, urban canyon, or densely populated environments [2].
To bridge this gap, 3GPP standardization bodies have foreseen, for the first time in 5th generation (5G) releases, the positioning service as a fundamental feature of the cellular network. Specifically, they introduced the Location Management Function (LMF) as one of the VNFs that can be instantiated within the Core Network (CN). However, the centralized architecture of the LMF introduces scalability and latency constraints, as the CN is typically deployed in a limited number of national-level sites. This centralization limits its applicability in scenarios demanding low-latency or high-precision positioning, particularly when rapid adaptation to dynamic radio conditions is required.
A more promising vision is to move positioning intelligence closer to the network edge, embedding it directly within the RAN. In this context, the architectural innovations introduced by the Open Radio Access Network (O-RAN) Alliance provide a natural enabler [3]. By exposing open interfaces and programmable control loops, said RAN Intelligent Controllers (RICs), O-RAN allows the development of localization micro-services (xApps) that can seamlessly integrate with the radio stack.
Within this framework, uplink Angle of Arrival (AoA) estimation based on the User Equipment (UE)’s Sounding Reference Signal (SRS) emerges as a particularly attractive primitive. This approach operates transparently to the UE, requiring no additional hardware or signaling overhead, and can seamlessly coexist with other network-native features. By confining the computation to the network edge, it enables a fast, lightweight, and easily deployable positioning service. Moreover, when combined with uplink UE ranging techniques [4], this solution supports single-anchor localization, eliminating the need for inter-5G base station (gNB) synchronization and orchestration, procedures that are typically complex and costly. Consequently, AoA serves as a lightweight yet powerful building block for network-native, single-anchor positioning.
Indeed, several studies (e.g., [5, 6]) have investigated this promising research direction, primarily through simulation-based evaluations. In contrast, only a few works have addressed the physical deployment and testing of AoA-based positioning, as pursued in this paper. For example, the testbeds presented in [7] and [8] demonstrate the feasibility of AoA estimation at the hardware level but are limited to minimal setups operating solely at the physical layer. We introduce the first fully open-source experimental framework that enables unrestricted and systematic testing of a complete 5G stack, namely, a fully operational 5G network with an integrated AoA service supporting both prototype and Commercial Off-the-Shelf (COTS) UE devices. This setup allows us to account for fundamental real-world non-idealities arising from hardware impairments and propagation environment characteristics. Consequently, the proposed testbed provides a flexible and reliable platform for reproducible and extensive performance evaluation of network-native positioning techniques.
In particular, by integrating a scenario-based ray tracer with a commercial channel emulator, we enable the evaluation of 5G positioning services across a wide range of scenarios under realistic and fully repeatable propagation conditions. In parallel, the specific challenge of AoA estimation is investigated through an extensive experimental campaign.
Main Contributions. This work provides the following key contributions. i) We present the first fully open-source experimental setup for evaluating positioning techniques in real 5G networks, supporting both prototype and COTS UEs. ii) We design and integrate in a full-stack 5G network an uplink AoA-based positioning function operating at the network edge and leveraging UE SRSs. iii) We combine a scenario-based ray tracer — Sionna RT [9] — with a commercial channel emulator — Keysight PROPSIM — to reproduce realistic and fully repeatable propagation conditions. iv) The proposed framework inherently captures hardware and environmental non-idealities, bridging the gap between simulation and practical deployment. v) An extensive experimental campaign validates the feasibility and performance of uplink AoA-based positioning in real 5G conditions.
Paper structure. We organize the rest of the paper as follows: Section II provides the fundamental background, Section III dives into the details of the AoA estimation method, Section IV describes the testbed we built for validating the method, Section V details our extensive experimental campaign, and last, Section VI summarizes the findings.
II Background
II-A 5G physical layer
The 5G New Radio (NR) physical layer is built on a scalable Orthogonal Frequency Division Multiplexing (OFDM)-based resource grid that adapts to diverse service requirements. Time is organized into radio frames of 10 ms, each divided into subframes, slots, Resource Blocks (RBs), and OFDM symbols. The Number of slots per subframe ($N_{slot}^{subframe,\mu}$) depends on the numerology index $\mu$, which further influences the subcarrier spacing, defined as $15\cdot 2^{\mu}$ kHz. Within this grid, user data and control signaling are jointly mapped to subframes, slots, and subcarriers, enabling flexible allocation of radio resources with adjustable trade-offs between signaling overhead and accuracy. This flexible numerology allows for balancing throughput, latency, and coverage requirements. Moreover, 5G supports advanced multi-antenna techniques such as MIMO and beamforming, providing the spatial resolution needed to extract AoA information.
II-B Sounding Reference Signal (SRS)
Among the various reference signals defined in 5G NR, the uplink SRS plays a central role in this work. Originally designed for channel sounding, uplink scheduling, and beam management, the SRS is a deterministic, unmodulated sequence transmitted by the UE and known at the gNB [10]. Its structure is based on Zadoff–Chu sequences, characterized by constant amplitude and zero autocorrelation properties, ensuring high robustness to interference and make them well suited for both channel estimation and positioning applications.
The SRS can be flexibly configured in both time and frequency domains. It may occupy one or more contiguous OFDM symbols within a slot and can be mapped across interleaved RBs over a configurable bandwidth. This flexibility enables the multiplexing of multiple UEs and promotes efficient reuse of radio resources. The network’s knowledge of the SRS sequence and its allocation in the resource grid makes it particularly suitable for uplink-based localization, allowing joint Time of Arrival (ToA) and AoA estimation without any hardware or software modifications to the UE. In this work, we exploit SRS transmissions to perform AoA estimation using subspace-based algorithms.
II-C Angle-of-Arrival (AoA) techniques
The potential of AoA-based techniques was first explored in the 1990s by [11] through the development of the Joint Angle and Delay Estimation (JADE) family of algorithms, which enable the joint extraction of AoA and Time of Flight (ToF). However, the integration of such techniques into mobile RANs has attracted significant attention only in the past decade. Notable contributions have focused on LTE systems [12], while the architectural innovations introduced with 5G RAN have further revitalized interest in this field. Despite their theoretical effectiveness, JADE-based algorithms are computationally demanding, mainly due to the eigendecomposition of large covariance matrices and the associated two-dimensional spectral search. Consequently, recent research efforts have shifted toward the adoption of more lightweight and robust AoA and ToA estimation methods leveraging native 5G signaling.
III AoA Estimation Framework
In this section, we present the framework used to estimate the AoAs of connected UEs from a serving gNB. In the proposed setup, a single anchor — namely, the gNB— is employed, equipped with an $M$-element uniform linear array (Uniform Linear Array (ULA)). However, the proposed method can be readily extended and integrated with ToF-based ranging techniques [4] to enable single-anchor localization.
III-A AoA estimation
To further assess the systematic capability of the testbed, we decided to test the performance of two different subspace-based techniques under the same environment and channel conditions. These techniques are based on the observation that the array covariance matrix can be decomposed into orthogonal signal and noise subspaces. The steering vectors corresponding to true AoA lie in the signal subspace, hence resulting in an orthogonal direction to the noise subspace.
To model the received signal, we assume a ULA of $M$ antenna elements, inter-spaced by $\Delta d=\lambda/2$, ensuring coverage over the range $90^{\circ},+90^{\circ}]$. Under the narrowband and far-field assumptions, which are reasonable for SRS and 5G RAN deployments, a source impinging from angle $\theta$ introduces a deterministic phase shift between adjacent antennas, captured by the steering vector
$\mathbf{a}(\theta)=\begin{bmatrix}1&e^{j\mu}&e^{j2\mu}&\cdots&e^{j(M-1)\mu}\end{bmatrix}^{T}$ (1)
where $\mu=-\dfrac{2\pi}{\lambda}\Delta d\cdot\sin(\theta)$ is the spatial frequency, and $T$ represents the transpose operator.
The ULA received signals at time $t$ are thus expressed as
$\mathbf{x}(t)=\mathbf{a}(\theta)s(t)+\mathbf{n}(t),$ (2)
where $s(t)$ denotes the transmitted signal after propagation attenuation, and $\mathbf{n}(t)$ represents an additive white Gaussian noise (AWGN) vector. In our case, $s(t)$ corresponds to the uplink SRS. Since SRS transmissions are orthogonal across users, the single-source model is valid, thereby simplifying the estimation process. For clarity, the following mathematical formulation focuses on the single-source case; however, the same derivations can be extended to multi-source AoA estimation when the number of sources is known.
The covariance matrix $\mathbf{R}_{x}$ captures the spatial correlation among the signals received by the $M$ antenna elements and is defined as
$\mathbf{R}_{x}=\mathbb{E}\{\mathbf{x}(t)\mathbf{x}^{H}(t)\}=\mathbf{a}(\theta)\mathbf{R}_{s}\mathbf{a}^{H}(\theta)+\sigma^{2}\mathbf{I},$ (3)
where $\mathbf{x}(t)$ is the received signal vector across the antenna array at time $t$, and $\mathbf{x}^{H}(t)$ denotes its Hermitian transpose. The operator $\mathbb{E}{\cdot}$ represents the statistical expectation, while $\mathbf{R}_{s}\triangleq\mathbb{E}{\mathbf{s}(t)\mathbf{s}^{H}(t)}$ is the source signal covariance matrix. The noise is assumed to be zero-mean Additive White Gaussian Noise (AGWN) with covariance $\sigma^{2}=\mathbb{E}{\mathbf{n}(t)\mathbf{n}^{H}(t)}$, and $\mathbf{I}$ denotes the identity matrix.
In practice, the covariance matrix is estimated from the phase-corrected received samples as
$\hat{\mathbf{R}}_{x}=\frac{1}{M_{sc,b}^{\text{SRS}}}\sum_{n=0}^{M_{sc,b}^{\text{SRS}}-1}\mathbf{x^{\prime}}(n)\mathbf{x^{\prime}}^{H}(n),$ (4)
where $M_{sc,b}^{\text{SRS}}$ denotes the number of SRS samples within the considered bandwidth, and $\mathbf{x}^{\prime}(n)$ represents the $n$-th phase-corrected SRS sample received at the ULA. By performing the eigendecomposition of $\hat{\mathbf{R}}_{x}\triangleq\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{H}$, we get $\boldsymbol{\Lambda}=diag\{\lambda_{0},\dots,\lambda_{M}\}$ containing the sorted eigenvalues, and $\mathbf{V}$ their associated eigenvectors, of which the leftmost M1M-1 correspond to the noise subspace, while the remaining ones define the signal subspace $\mathbf{V}_{s}$.
III-A1 MUSIC
The MUltiple SIgnal Classification algorithm [13] offers higher resolution compared to classical beamforming techniques [14], as it enables the discrimination of closely spaced multipath components even under limited Signal-to-Noise-Ratio (SNR) conditions.
The algorithm exploits the orthogonality condition between the signal and noise subspaces, expressed as
$\mathbf{a}^{H}(\theta)\mathbf{V}_{n}\mathbf{V}_{n}^{H}\mathbf{a}(\theta)=0,$ (5)
where $\mathbf{V}n$ denotes the matrix of noise eigenvectors, i.e. the noise subspace of $\hat{\mathbf{R}}_{x}$. This condition leads to the MUSIC pseudospectrum $P(\theta)$, whose peaks correspond to the estimated angles of arrival $\hat{\theta}$:
$\hat{\theta}=\arg\max_{\theta}\{P(\theta)\}\\ =\arg\max_{\theta}\Bigl\{\frac{1}{\mathbf{a}^{H}(\theta)\mathbf{V}_{n}\mathbf{V}_{n}^{H}\mathbf{a}(\theta)}\Bigr\}.$ (6)
In the proposed uplink SRS-based configuration, this approach provides high-resolution user direction estimation with moderate computational complexity, making it suitable for potential real-time deployment.
III-A2 ESPRIT
The Estimation of Signal Parameters via Rotational Invariance Techniques overcomes MUSIC’s computational and storage requirements by exploiting the shift-invariance property of the array [15].
ESPRIT assumes that the ULA can be decomposed into two (possibly overlapping) equally-sized sub-arrays, $sub_{0}$ and $sub_{1}$. The received signal at each $sub_{i}$ is modeled as in (2):
$\begin{split}\begin{bmatrix}\mathbf{x}_{0}(t)\\ \mathbf{x}_{1}(t)\end{bmatrix}=\begin{bmatrix}\mathbf{a}(\theta)\\ \mathbf{a}(\theta)\mathbf{\Phi}\end{bmatrix}\mathbf{s}(t)+\begin{bmatrix}\mathbf{n}_{0}(t)\\ \mathbf{n}_{1}(t)\end{bmatrix}=\tilde{\mathbf{a}}(\theta)\mathbf{s}(t)+\mathbf{n}(t).\end{split}$ (7)
where $\mathbf{a}(\theta)$ is the steering vector as in (1). The signal received by $sub_{1}$ will experience an extra delay due to the fixed displacement $\Delta d$ among antennas, described by the spatial rotational matrix $\mathbf{\Phi}=[e^{j\mu}]$ (a diagonal matrix for multi-source). The objective of the ESPRIT technique is to estimate the AoA by estimating $\mu$ through determining $\mathbf{\Phi}$. In doing so, two steps are required: estimate the signal subspace $\mathbf{V}_{s}$ derived from $\hat{\mathbf{R}}_{x}$ as seen in (4), and then estimate the spatial rotational matrix $\mathbf{\Phi}$.
Once the estimated signal subspace $\mathbf{V}_{s}$ is obtained, it can be decomposed into $\mathbf{V}_{s0}$ and $\mathbf{V}_{s1}$ for $sub_{0}$ and $sub_{1}$ respectively, to form the so-called invariance equation $\mathbf{V}_{s0}\mathbf{\Psi}=\mathbf{V}_{s1}$, from which we can estimate the subspace rotating operator $\hat{\mathbf{\Psi}}$ by means of least-squares technique. Due to the shift-invariance property, it holds that
$\hat{\mathbf{\Psi}}=\mathbf{T}\mathbf{\Phi}\mathbf{T}^{-1},\;\ \exists\,\mathbf{T}\ \text{s.t.}\ (\det\mathbf{T}\neq 0\;\wedge\;\mathbf{a}(\theta)=V_{s}\mathbf{T}),$ (8)
from which we calculate the eigenvalues of the resulting complex-valued solution, and then extract the AoA information from $\mathbf{\Phi}$ via
$\hat{\theta}=\arcsin\biggl(-\frac{\lambda}{2\pi\Delta d}\mu\biggr).$ (9)
III-B Calibration
Figure 1: over-the-air calibration example, by using a 4-element ULA.
One of the main challenges in real-world AoA estimation is the hardware-dependent phase misalignment among antenna ports. In our setup, the employed Software-defined Radio (SDR) is equipped with two separate daughterboards, each featuring an independent Radio Frequency (RF) chain and local oscillator. As stated in the manufacturer’s documentation:
“The N310 has no way of aligning the phase between channels, and it will be random between runs.” [16]
As a result, the relative carrier phases are not inherently synchronized at startup, introducing an arbitrary phase offset between antenna ports. Without proper correction, this offset prevents accurate AoA estimation by algorithms that rely on inter-antenna phase differences.
To ensure that the inter-antenna phase difference remains stable for the duration of each measurement campaign, we modeled a first over-the-air (OTA) calibration mechanism, performed opportunistically at system start-up.
A reference UE is temporarily positioned at $0^{\circ}$ in front of the array, where the theoretical inter-antenna phase difference is zero under narrowband and far-field assumptions, conditions satisfied by the transmitted SRS and the UE placement. From the received uplink SRS sequences, the instantaneous phase offsets of each antenna port relative to the reference port 0, denoted $\Delta\hat{\phi}_{0,m}$, where $m\in(0,M]$, are estimated and averaged over a short time interval to obtain $\Delta\bar{\phi}_{0,m}$. These offsets represent per-port phase biases that remain constant until the next SDR reboot and are stored for the duration of the measurement campaign.
During test operations, each received sample $x_{m}(n)$ from port $m$ is thus digitally corrected as
$x_{m}^{\prime}(n)=x_{m}(n)\cdot e^{-j\Delta\bar{\phi}_{0,m}}\quad for\;n\in[0,M_{sc,b}^{\text{SRS}}-1],$ (10)
where $M_{sc,b}^{\text{SRS}}$ denotes the length of the SRS sequence, corresponding to $960$ samples in the considered full-bandwidth configuration at 6060MHz. This procedure ensures consistent phase alignment across antennas during multiple UE attachment and detachment cycles.
This method requires no external reference or additional hardware, embedding the calibration into normal network operation and ensuring stable inter-antenna phase coherence for AoA estimation.
III-C Cylindrical Correction
Figure 2: Geometric scenario of cylindrical correction. MUSIC naturally estimates the orange angle.
An additional correction is applied to ensure that the positioning framework operates on the $XY$ plane, estimating the top-view azimuthal angle, green in Fig. 2, rather than on the plane that intersects the ULA and the UE, which corresponds to the raw AoA (orange in Fig. 2).
The correction formula can be expressed as:
$\hat{\theta}_{XY}=\arcsin{\biggl(\frac{d\sin\hat{\theta}}{\sqrt{d^{2}-(\Delta z)^{2}}}\biggr)},$ (11)
where $d$ denotes the ULA-UE distance, which can be obtained by integrating ranging information from a complementary positioning service, and $\Delta z$ represents the vertical offset between the ULA and the UE. The height of the ULA is assumed to be known for each scenario, while the UE height is fixed at $1.5$ m.
Figure 3: Experimental testbed: at the top, the Sionna ray tracer, which fetch the channels to the PROPSIM; to the left, the UE connected through a douplex connection to the O-RAN gNB.
IV Experimental Setup
To validate the proposed framework, we developed an experimental setup enabling a comprehensive evaluation of the integrated AoA microservice across diverse propagation scenarios. The setup allows fine-tuning of channel characteristics to accurately reproduce realistic radio conditions.
Ray-tracing scenarios were designed in Sionna RT  [^1] https://developer.nvidia.com/sionna for channel characterization in specific environments. Each scenario includes a line-of-sight (LOS) component, along with specular and diffuse surface reflections. As the experiments were conducted at 3.95GHz, refraction effects were considered negligible. For each scenario, we modeled LOS as well as third- and fifth-order reflections to assess the impact of multipath propagation on AoA estimation accuracy.
In all virtual scenarios, the gNB was positioned at fixed coordinates $(x_{\textit{gNB}},y_{\textit{gNB}},z_{\textit{gNB}})$ for the entire duration of the experiments. It was equipped with a four-element isotropic antenna array arranged in a $1\times 4$ planar configuration, with elements spaced by $\frac{1}{2}\lambda$. Conversely, the UE employed a single vertically polarized isotropic transmitting element. The UE was modeled in Sionna as a moving object following a trajectory specified in an external .csv file , which provides its coordinates at discrete time steps $p_{\textit{UE}}(k)=k\cdot\Delta t$ for $k\in[0,900]$, with $\Delta t=0.1$ s. For each transmitter position, a path solver instance computed all UE-gNB propagation paths up to the specified interaction depth (i.e., reflection order). This process generated a set of multi-antenna, time-varying Channel Impulse Responses (CIRs) describing the propagation characteristics of a dynamic urban environment.
We then used Keysight PROPSIM F8800A channel emulator, a device capable of reproducing the exact propagation conditions of a target environment in real time, while remaining transparent to the connected devices. Channel emulation enables interaction with actual hardware and software as if they were operating in their native radio context, making it particularly valuable for testing scenarios that would otherwise be difficult, costly, or time-consuming to realize in practice. This setup allows us to combine the flexibility of simulation-based configuration with the realism of hardware-based validation. The CIRs generated in Sionna RT are converted into the PROPSIM-compatible .asc format using a dedicated channel converter developed for this framework.
For the validation phase, we integrated both commercial devices and devices based on open-source software, as depicted in Fig. 3. The UE is a COTS IoT device, a Sierra Wireless unit mounting a Qualcomm Snapdragon 888 5G modem. Its transceiver is connected by coaxial cables to a port of the Keysight PROPSIM, which emulates the propagation channels. On the gNB side, one emulator port is reserved for downlink traffic, while four ports are for uplink transmissions, corresponding to the four antenna elements of the Radio Unit (RU). The RU is implemented using an Ettus Research N310 SDR equipped with two daughterboards, each providing two transmit and two receive antenna ports. The SDR streams digitized baseband data to the Distributed Unit (DU) through a dedicated Network Interface Card using two 10 Gb/s Small Form-factor Pluggable (SFP) + optical links. The gNB is deployed on a general-purpose server running the open-source OAI software stack and is connected to the Open5GS open-source 5G core network.
The programmability of the open RAN software stack enables the seamless integration of our dedicated solution within the OAI framework, allowing controlled experimentation with commercial devices and open-source network components under emulated yet realistic channel conditions.
V Experimental Results
This section presents a systematic assessment of the proposed uplink AoA estimation framework using the testbed presented in Sec. IV.
The experiments were conducted across the aforementioned three scenarios designed to reproduce one of the most challenging geometries for wireless communication systems: urban canyons, which are known to degrade localization performance due to their multipath-rich propagation conditions. We evaluated the accuracy of the proposed framework over a total of 69,303 SRS signals exchanged between UE and gNB, $6\%$ of which correspond to Non-Line-of-Sight (NLOS) conditions. Among all the signals exchanged, $31\%$ were in a zero multipath scenario, $34\%$ in a 3rd-order multipath environment, and $35\%$ in a 5th-order one.
V-A The impact of multipath
(a) No multipath scenario, LOS only.
(b) Up to 5-th order reflections.
Figure 4: An example of AoA estimation with and without multipath for the same urban scenario. In orange the estimated $\hat{\theta}$, while in green, the corrected $\hat{\theta}_{XY}$.
By progressively increasing the maximum reflection order, the channel is enriched with additional reflected components, allowing the analysis of the framework’s performance under increasingly complex and challenging propagation conditions.
Fig. 4 illustrates a representative AoA estimation trace obtained with (Fig. 4(b)) and without (Fig. 4(a)) multipath propagation. As the maximum reflection order increases, the number of secondary paths grows, leading to higher estimation variance and shifts in the MUltiple SIgnal Classification (MUSIC) pseudospectrum. This behavior is evident from the numerous spikes observed in Fig. 4(b), which, however, can be effectively smoothed using a Kalman filter, as demonstrated in [17, 18]. This effect becomes particularly pronounced in NLOS conditions, highlighted in gray in Fig. 4. Although multipath propagation reduces the accuracy of AoA estimation, it remains crucial for maintaining UEgNB connectivity when the direct path is obstructed. As shown in Fig. 4(a), the UE disconnects from the gNB when both the line-of-sight component and sufficient power from reflected paths are absent, resulting in link failure.
Figure 5: MUSIC vs ESPRIT in three different multipath depth scenarios, after cylindrical correction.
We further assess the impact of multipath using Fig. 5, which depicts the Empirical Cumulative Distribution Function (ECDF) of the Root Mean Squared Error (RMSE) for different reflection orders. The results confirm that in the absence of multipath (solid curves), the estimation error is tightly concentrated, with over 95% of realizations exhibiting an RMSE below $2^{\circ}$. As the channel becomes progressively enriched with reflected components (dashed and dotted curves), the error distribution widens, revealing the degradation introduced by secondary propagation paths. Overall, the ECDF analysis corroborates that although multipath is essential to sustain connectivity under NLOS conditions, it also constitutes the main source of estimation variability, particularly at higher reflection orders. Consequently, as expected, MUSIC and ESPRIT exhibit comparable accuracy under the same environmental conditions.
V-B The impact of SNR
Figure 6: Distribution of RMSE Pre- vs Post- plane correction across different SNR ranges with MUSIC and ESPRIT AoA estimation.
The accuracy of the proposed framework improves as the SNR increases. Although this trend is expected, it is important to quantify the SNR range over which the AoA service remains effective. The boxplot in Fig. 6 illustrates the distribution of AoA RMSE across different SNR ranges. An improvement in bias-corrected estimation accuracy is observed as the SNR increases: for SNR values above 30 dB, over 95% of the estimates exhibit an RMSE below $3^{\circ}$, while in the 0–10 dB range, achieving the same 95% level requires tolerating errors up to $9^{\circ}$. These results demonstrate that SNR is the dominant factor influencing AoA estimation accuracy. However, even in multipath-rich environments, when the SNR is sufficiently high, the MUSIC algorithm consistently achieves sub-degree to few-degree precision. Conversely, under low-SNR conditions, noise degrades performance and increases variability; nevertheless, the algorithm continues to provide estimates that remain within practical accuracy bounds.
VI Conclusions and Future Work
In this work, we presented a flexible and controllable testbed that, for the first time, enables systematic evaluation of AoA-based positioning techniques under real yet fully repeatable channel conditions. The proposed framework integrates NVIDIA Sionna RT with the Keysight PROPSIM channel emulator and includes a novel phase calibration procedure for the USRP N310. By combining simulation flexibility with hardware realism, the testbed allows the experimental validation of AoA estimation algorithms in realistic propagation environments, providing results of direct relevance for next-generation 5G deployments.
The proposed opportunistic calibration method proved effective but still requires manual initialization. Future work will focus on automating this process through self-calibration mechanisms that exploit known pilot signals transmitted and analyzed by the gNB itself. Such automation would reduce operator intervention and ensure long-term phase stability, enabling fully autonomous operation of the testbed.
Finally, the AoA estimation algorithm has been implemented for deployment as an O-RAN xApp, enabling seamless integration with complementary ToA-based frameworks. This makes the proposed platform a practical and extensible foundation for network-native 5G localization services.